# Mining the Minoria: Unknown, Under-represented, and Under-performing Minority Groups

### v1

@@ 6.3. Focused Exploration
A significant challenge for the E&E approach is that in high-dimensional cases where $d\gg 2$, the search space becomes huge, making it unlikely to find near-optimal directions. Therefore, instead, we first identify the under-performing regions with the hope of reducing the search space significantly. To do so, we use the model $h_{\theta}$, trained on $\mathcal{D}$, to find a subset $S\subset\mathcal{D}$, which we call the error region, for which $h_{\theta}$ under-performs. Formally, the expected loss of the model in $S$ is significantly higher compared to $\mathcal{D}\setminus S$.
@@
Many implementation details are deferred to § 7 for ease of explanation. A simplified pseudo-code of this approach is given in Algorithm  3. Based on our experiments, this approach works well in practice on large-scale high-dimensional settings.
@@ 7. Experiments
Having presented our algorithms in this section, we evaluate them experimentally for the 2D case and higher dimensions. We analyze our algorithms in each section and their effectiveness in finding the high-skew directions and candidate Minoria. The algorithms are implemented using Python, and the experiments are conducted on a workstation with Ubuntu 18.0 OS. The codes are available at https://github.com/UIC-InDeXLab/Mining_U3Ms. We used k-means clustering as a baseline in § 7.1.2.
@@
(a) First high-skew direction.
(b) Second high-skew direction.
Figure 3. Top-$2$ high-skew directions on the Latitude and Longitude of Chicago Crimes dataset are presented in these two plots (the plots show a smaller sample of the dataset points). In each plot, the output of the binary classification model is presented with green and blue lines. The red line shows the direction $f$ with the highest projection skew. The orange region shows the tail region.
@@ 7.1.1. Chicago Crimes
Table 2. (Chicago crime) Performance evaluation on second high-skew direction (skew = 0.16). The tail is calculated on the entire dataset points.
Results
The top-$2$ high-skew directions are shown in Fig 3. As we can see, the first two high-skew directions are along the north side to the south side of Chicago. We select $0.01$, $0.001$, and $0.0001$ percentile of the points in the tail of the high-skew projected points. The tails of the high-skew directions are presented using the orange regions. Evidently, in both cases, the tails correspond to the North-side Chicago. To provide some background about the city, Chicago is one of the most segregated cities in the US, with the vast majority of its north side being white, while the south side is black majority. We refer the curious reader to [17, 28] for more details. Interestingly, the high-skew projections in this dataset discovered a valid Minoria– representing the under-represented and under-performing group of white people. The white group is under-represented in the crime dataset because of many reasons, such as predictive policing [15]. At the same time, the f1-score significantly decreases on the tail of the projection, illustrating that predictions based on the crime dataset are not accurate for this group. To provide more details about these directions, we provide the tail-performance for different percentiles for these two directions in Tables 1 and 2. As illustrated in the tables, although the accuracy shows only a slight variation, there is a significant decrease in the F1-score on the tails in both cases. This indicates a pronounced class imbalance in the tail distribution, which is attributable to the demographic group imbalance.
@@ Results
In Fig  4(a), we provide a visualization of GRE and GPA values from this dataset. The red line shows the highest skew direction. Fig  4(b) visualizes the tail of this high-skew direction (percentile $p=0.1$). As we can see in Table  4, the ratio of the Female group is significantly larger in the $0.1$, $0.08$, and $0.04$ percentiles compared to the whole dataset (1.1). The accuracy decreases up to %10 percent when concentrating on the tail of this direction. In this case, the Female group is not under-represented in the dataset because the ratio is almost 1.0. However, it is an under-performing and unknown minority group.
To compare our proposed method with existing clustering techniques, we applied k-means clustering with $k=5$ to the same dataset. The clusters are visualized in fig  4(c). The female-to-male ratio is calculated in Table  4. Most clusters do not appear suitable candidates for detecting underrepresented demographic groups. This is because the ratio of females to males within these clusters differs from the overall dataset ratio (1.1) by at most 0.4. However, applying the ray-sweeping algorithm enabled the identification of the $0.1$-tail, which contains a comparable number of points to Cluster 1 but exhibits a significantly higher female-to-male ratio. This suggests that the tail may serve as a more suitable candidate than the best cluster identified by the k-means algorithm. Increasing the value of $k$ in the k-means clustering would result in more clusters to evaluate, thereby extending the computational time required. However, reducing this number leads to fewer clusters, each containing a larger number of points, which makes it more challenging for a human expert to discern patterns within them. This analysis indicates that our method demonstrates that clustering failed to identify the underrepresented minorities.
@@
Percentile Accuracy F1-score Female/Male ratio on Tail
@@
Cluster id Size Female/Male ratio
0 145 1.11
1 45 1.49
2 72 0.95
3 83 0.24
@@
Table 4. Evaluation of the output clusters of k-means algorithm in detecting the unknown minority groups.
@@ 7.2. Higher-dimensional Experiments
In this section, we evaluate the higher-dimensional methods proposed in § 6. As reflected in our negative result on the error-bound of the discretization approaches ( § 6.1.3), their effectiveness depends on the dataset topology. Besides, their running time significantly increases by the number of dimensions. Our experiments verify these findings: we found discretization slow and less effective. Therefore, in the following we focus on the more practical approaches based on the search space exploration. Comparing the E&E method (§ 6.2) and the focused exploration approach (§ 6.3), our experiments found the latter (focused exploration) significantly more effective for Minoria mining. Therefore, in the following, we first present our results for this approach and then present our extended experiment results for the E&E approach and highlight some of its drawbacks.
@@ Results
In these experiments, we hide the sensitive attribute containing the under-performing minority group and check if our mining experiments can discover it. To do so, we solved Problem 1 while setting $\ell=5$. For the Adults and Diabetes datasets, we considered female as the minor demographic group. For the COMPAS dataset, we considered the race Caucasian (white) as the minor demographic group. The result of running the algorithm on the three datasets is shown in Fig  5. For the Adults dataset in Fig  5(a), three of the five high-skew directions successfully show the Female group as the Minoria. This is because when focusing on the high error points in the tail of these directions, the ratio of Female points increases significantly. In addition, we can see that the ratio of the Female group in the tail of these high-skew directions is larger than this ratio for the whole dataset. This shows that our model did not perform well on the minor group. For the COMPAS dataset in Fig  5(b), two of the high-skew directions show the Race Caucasian as the Minoria. The red line is the high-skew direction that shows the model is significantly under-performing on this minor group. For the Diabetes dataset  5(c), all the high-skew directions show that we have more error in the regions where the ratio of the minor group is larger than this ratio in the whole dataset. However, it is not easy for a human expert to detect this minor group as problematic because the ratio of this group is almost less than 0.3 in the high error subset of the tail. For the Diabetes dataset (Fig  5(c)), the tail skewed directions tend to show the problematic points as minor groups. As we focus on the top high error points of the tail of each direction, the ratio of the minor group increases. However, since this ratio is not large enough (less than 0.5), showing the points to a human supervisor may not determine the exact problematic feature in this region.
@@
Figure 5. The result of running the Focused Exploration method on the three datasets. Each line corresponds to one of the top-$5$ high-skew directions. The brown line shows the ratio of minor group in the whole dataset. The x-axis shows different values for a hyper-parameter $k$. The ratio of minorities (y-axis) is calculated on the top-$k$ high error points in the tail of each high-skew direction. If the ratio increases while decreasing $k$, we have more of the minor group in the set of high error points of the tail of the specific direction.
@@
First, we explored a case when the demographic groups are well-separated linearly. This means there can be a linear classifier to classify these groups with an accuracy of more than %80. In this case, the E&E algorithm successfully finds the high-skew directions that separate the minor group from the major one. In Fig  6, the top-$3$ high-skew directions are found by this method. The plots show the histogram of the projected points on the high-skew direction $f$. In this case, by picking each direction’s tail, we successfully target the minor group (red points). However, consider a case when the two demographic groups are hard to separate linearly (Fig  7). The top-$3$ high-skew directions do not help find the demographic groups in this case. The red points are almost overlapped with the major blue points. As a result, reporting the tail of this skew to a human expert does not help him to detect a Minoria.
@@
Figure 6. An illustration of the projected points on the three highest skewed directions discovered by the E&E method. In this case, the demographic groups (hidden during the learning phase) are well separated in the original dataset (the accuracy of a Logistic Regression classifier is almost %80). The red points belong to the minor group.
Figure 7. An illustration of the projected points on the three highest skewed directions discovered by the E&E method. In this case, the demographic groups (which are hidden during the learning phase) are not linearly separable in the dataset space. The red points belong to the minor group.
@@ 9. Conclusion
In this paper, we introduced, formalized, and studied the problem of mining the unknown, underrepresented, and underperforming minority groups (Minoria). Making the connection to computational geometry concepts, we proposed efficient algorithms and practical solutions for finding high-skew projections as the combination of data features that reflect high performance disparities at their tail. The scope of our proposed techniques is limited to the cases where the Minoria are linearly separable from the rest of the data. Extending our work to the more general class of non-linear projections, as well as developing alternative Minoria mining approaches is an interesting direction for future work.

### v2

@@ 6.3. Focused Exploration
A significant challenge for the E&E approach is that in high dimensional cases where $d\gg 2$, the search space becomes huge, making it unlikely to find near-optimal directions. Therefore, instead, we first identify the under-performing regions with the hope of reducing the search space significantly. To do so, we use the model $h_{\theta}$, trained on $\mathcal{D}$, to find a subset $S\subset\mathcal{D}$, which we call the error region, for which $h_{\theta}$ under-performs. Formally, the expected loss of the model in $S$ is significantly higher compared to $\mathcal{D}\setminus S$.
@@
Many implementation details are deferred to § 7 for ease of explanation. A simplified pseudo-code of this approach is given in Algorithm  2. Based on our experiments, this approach works well in practice on large-scale high-dimensional settings.
@@ 7. Experiments
Having presented our algorithms in this section, we evaluate them experimentally for the 2D case and higher dimensions. We analyze our algorithms in each section and their effectiveness in finding the high-skew directions and candidate Minoria. The algorithms are implemented using Python, and the experiments are conducted on a workstation with Ubuntu 18.0 OS.
@@ 7.1.1. Chicago Crimes
(a) First high-skew direction.
(b) First high-skew direction. After projecting all the points.
Figure 4. Top high-skew directions on the Latitude and Longitude of Chicago Crimes dataset are presented in these two plots (the plots show a smaller sample of the dataset points). In the first plot, the output of the binary classification model is presented with green and blue lines. The red line shows the direction $f$ with the highest projection skew. The orange region shows the tail region.
Results
The top-$2$ high-skew directions are shown in Fig 4. As we can see, the first two high-skew directions are along the north side to the south side of Chicago. We select $0.01$, $0.001$, and $0.0001$ percentile of the points in the tail of the high-skew projected points. The tails of the high-skew directions are presented using the orange regions. Evidently, in both cases, the tails correspond to the North-side Chicago. To provide some background about the city, Chicago is one of the most segregated cities in the US, with the vast majority of its north side being White  [^8] Please note that north-side being a Minoria (i.e., under-represented and under-performing) in the Crimes dataset is irrespective of its segregated nature.. We refer the curious reader to [18, 30] for more details. North-side is under-represented in the dataset for many reasons, such as predictive policing [15]. At the same time, the f1-score significantly decreases on the tail of the projection, illustrating that predictions based on the crime dataset are not accurate for this group. To provide more details about these directions, we provide the tail-performance for different percentiles for these two directions in Tables 1 and 2. As illustrated in the tables, although the accuracy shows only a slight variation, there is a significant decrease in the F1-score on the tails in both cases. This indicates a pronounced class imbalance in the tail distribution, which is attributable to the demographic group imbalance.
@@ Results
Table 2. (Chicago crime) Performance evaluation on second high-skew direction (skew = 0.16). The tail is calculated on the entire dataset.
@@
In Fig  5(a), we provide a visualization of GRE and GPA values from this dataset. The red line shows the highest skew direction. Fig  5(b) visualizes the tail of this high-skew direction (percentile $p=0.1$). As we can see in Table  5, the ratio of the Female group is significantly larger in the $0.1$, $0.08$, and $0.04$ percentiles compared to the whole dataset (1.1). The accuracy decreases up to %10 percent when concentrating on the tail of this direction. In this case, the Female group is not under-represented in the dataset because the ratio is almost 1.0. However, it is an under-performing and unknown minority group.
Runtime
The running time of the Ray-sweeping algorithm on the Chicago Crime dataset in our settings was $8$ seconds. This time was $3.74$ seconds for College-admission dataset.
7.1.3. Comparison with Baselines
We utilized two clustering techniques as baselines: k-means [28] and Affinity Propagation [23]. The resulting clusters were analyzed to identify potential candidates for the underrepresented demographic groups (Minoria). We employed k-means with various values of $k$ and utilized Affinity Propagation with a carefully adjusted preference parameter to control the number of clusters and avoid excessive clustering. A visualization of the k-means clusters on College-admission dataset is provided in Figure 5(c).
The ratio of minority groups within each cluster in k-means for various datasets is presented in Table 5. As reflected in the table, unlike the region discovered using our methods (tail of skews), the ratios of minority groups within each cluster are not significantly different from the whole dataset (last row). As a result, these clusters are not good candidates for identifying underrepresented groups within the data. Furthermore, the number of points in each cluster is significantly larger compared to the regions we identified, making it more challenging to detect and analyze minority groups within such clusters. In contrast, the regions we discovered are more compact and therefore more efficient for mining underrepresented groups, offering a practical and focused approach to this task. Table 5 compares the results for various values of $k$, all yielding consistent results.
The results of applying Affinity Propagation clustering are presented in Figure 7(b). Similar challenges persist, including an excessive number of clusters, requiring all clusters to be examined. Furthermore, the cluster ratios do not significantly differ from those observed in the entire dataset.
Runtime
The average run-time of k-means on the College admission dataset in our setting is $0.003$ milliseconds. This value for the Affinity Propagation method is $0.13$ seconds. Even though the running time is smaller for clustering methods compared to the Ray-sweeping algorithm, their output clusters are not valid for solving the Minoria Mining problem defined in this work.
@@ 7.1.3. Comparison with Baselines
Percen- Accu- F1 Female/Male
tile racy score ratio on Tail
@@
Cluster Female/Male
ID Size ratio
@@
1 72 0.94
2 108 1.11
3 45 1.50
4 83 1.24
Total 400 1.10
Table 4. Evaluation of the output clusters of k-means clustering in detecting the unknown minority groups on College-admission dataset with $k=5$.
$k$ Max Size Maximum F/M ratio
2 201 1.26
3 136 1.20
4 104 1.20
5 108 1.50
6 90 1.44
7 93 1.50
Total 400 1.10
Table 5. Varying $k$ in k-means on College-admission dataset. Col 2 shows the max Female to Male ratio on the output clusters.
@@ 7.2. Higher-dimensional Experiments
In this section, we evaluate the higher-dimensional methods proposed in § 6. As reflected in our negative result on the error-bound of the discretization approaches [16], their effectiveness depends on the dataset topology. Besides, their running time significantly increases by the number of dimensions. Our experiments verify these findings: we found discretization slow and less effective. Therefore, in the following we focus on the more practical approaches based on the search space exploration. Comparing the E&E method (§ 6.2) and the focused exploration approach (§ 6.3), our experiments found the latter (focused exploration) significantly more effective for Minoria mining. Therefore, in the following, we first present our results for this approach and then present our extended experiment results for the E&E approach and highlight some of its drawbacks.
@@ Results
In these experiments, we hide the sensitive attribute containing the under-performing minority group and check if our mining experiments can discover it. To do so, we solved Problem 1 while setting $\ell=5$. For the Adults and Diabetes datasets, we considered female as the minor demographic group. For the COMPAS dataset, we considered the race Caucasian (white) as the minor demographic group. The result of running the algorithm on the three datasets is shown in Fig  6. For the Adults dataset in Fig  6(a), three of the five high-skew directions successfully show the Female group as the Minoria. This is because when focusing on the high error points in the tail of these directions, the ratio of Female points increases significantly. In addition, we can see that the ratio of the Female group in the tail of these high-skew directions is larger than this ratio for the whole dataset. This shows that our model did not perform well on the minor group. For the COMPAS dataset in Fig  6(b), two of the high-skew directions show the Race Caucasian as the Minoria. The red line is the high-skew direction that shows the model is significantly under-performing on this minor group. For the Diabetes dataset  6(c), all the high-skew directions show that we have more error in the regions where the ratio of the minor group is larger than this ratio in the whole dataset. However, it is not easy for a human expert to detect this minor group as problematic because the ratio of this group is almost less than 0.3 in the high error subset of the tail. For the Diabetes dataset (Fig  6(c)), the tail skewed directions tend to show the problematic points as minor groups. As we focus on the top high error points of the tail of each direction, the ratio of the minor group increases. However, since this ratio is not large enough (less than 0.5), showing the points to a human supervisor may not determine the exact problematic feature in this region.
Runtime
The running time of Focused Exploration in our settings was $112$ seconds on Adults dataset, $24$ seconds on COMPAS, and $213$ seconds on Diabetes dataset.
7.2.2. Comparison with Baseline
In this subsection, similar to subsection 7.1.3, we compare the results with two clustering algorithms: k-means and Affinity Propagation. Tables 9 and 9 show the ratio of minorities on the output clusters of k-means with $k=5$. Same results as Section 7.1.3 is observed here. The result of comparing different values $k$ is shown in Tables 9 and 9. Applying Affinity Propagation will also result in clusters which are not good candidates for Minoria mining (Figure 7(a)).
Runtime
The running time of k-means clustering on Adults and COMPAS datasets was $0.12$ and $0.33$ seconds, respectively. This running time for Affinity Propagation method was $14$ and $20.2$ seconds on Adults and COMPAS datasets. Even though the running time is smaller for clustering methods, their output clusters are not valid for solving the Minoria Mining problem defined in this work.
@@ 7.2.2. Comparison with Baseline
Figure 6. Focused Exploration on the three datasets. Each line corresponds to one of the top-$5$ high-skew directions. The brown line shows the ratio of minor group in the whole dataset. The x-axis shows different values for a hyper-parameter $p$ (or $p$-percentile of the tail). The ratio of minorities (y-axis) is calculated on the top-$k$ (i.e. top $p$-percentile) high error points in the tail of each high-skew direction. If the ratio increases while decreasing $p$, we have more points from the minor group in the set of high error points of the tail of the specific direction.
Cluster Size female
ID ratio
0 10850 0.13
1 5718 0.30
2 14678 0.52
3 3952 0.47
4 13644 0.24
Total 48842 0.33
Table 6. Evaluation of the output clusters of k-means when $k=5$ algorithm in detecting the unknown minority groups for Adults dataset.
$k$ Max Max female
Size ratio
2 45252 0.33
3 19552 0.53
4 19522 0.52
5 14678 0.52
6 13041 0.53
7 13670 0.84
Total 48842 0.33
Table 7. Varying $k$ in k-means for Adults dataset. Col 2 shows the max Female/total ratio in output clusters.
Cluster Size Caucas-
ID ian ratio
0 524 0.09
1 5174 0.38
2 5860 0.29
3 2781 0.44
4 2680 0.25
Total 18316 0.33
Table 8. Evaluation of output clusters of k-means when $k=5$ algorithm in detecting the unknown minority groups for COMPAS dataset.
$k$ Max Max Cauc-
Size asian ratio
2 9020 0.35
3 7169 0.35
4 6638 0.45
5 5860 0.44
6 5100 0.46
7 4511 0.46
Total 18316 0.33
Table 9. Varying $k$ in k-means for COMPAS dataset. Col 2 shows the max Caucasian/total ratio in output clusters.
(a) Adults dataset (48,842 points)
(b) College admission dataset (400 ponints)
Figure 7. Affinity Propagation clustering: the size and the ratio of the minority group for each cluster are shown in different $y$-axices.
@@ Results
First, we explored a case when the demographic groups are well-separated linearly. This means there can be a linear classifier to classify these groups with an accuracy of more than %80. In this case, the E&E algorithm successfully finds the high-skew directions that separate the minor group from the major one. In Fig  8, the top-$3$ high-skew directions are found by this method. The plots show the histogram of the projected points on the high-skew direction $f$. In this case, by picking each direction’s tail, we successfully target the minor group (red points). However, consider a case when the two demographic groups are hard to separate linearly (Fig  9). The top-$3$ high-skew directions do not help find the demographic groups in this case. The red points are almost overlapped with the major blue points. As a result, reporting the tail of this skew to a human expert does not help him to detect a Minoria.
@@
Figure 8. Projected points on highest skewed directions discovered by E&E. In this case, the demographic groups (hidden during the learning phase) are well separated in the original dataset (the accuracy is almost %80). The red points belong to the minor group.
Figure 9. Projected points on highest skewed directions discovered by E&E. In this case, the demographic groups (hidden during the learning phase) are not linearly separable. The red points belong to the minor group.
@@ 8. Related Work
9. Discussion
In this section, we discuss some aspects of our proposed approach and highlight its limitations.
In general, our methods is limited to groups that are (almost) linearly separable. If the groups are linearly separable, the high-skewed direction found by our algorithms corresponds to a linear classification. In cases where linear separation is not feasible, one can consider introducing new features as polynomial combinations of existing features.
Not all the group candidates identified as the under-represented and under-performing regions in the data space are socially valid. Discovering such regions is intended to help model developers and data owners identify potential issues, decide if those are socially concerning, and proactively limit the scope of use of their datasets.
Our approach is only one of the possible solutions for Minoria mining. Finding alternative approaches requires extensive future research. Still, our experiments verified that clustering failed to identify the under-represented and under-represented groups.
Selecting appropriate columns from a dataset to perform Ray-sweeping is a critical practical decision. For instance, in our experiments, we utilized the Latitude and Longitude columns from the Chicago Crime dataset and the (GRE, GRA) pair from the College Admission dataset. One potential direction for column selection is identifying the most influential features for black-box machine learning models. Techniques such as feature importance measures or Shapley values can be employed to score each column, allowing the selection of the most impactful features for this algorithm.
In higher dimensions, the curse of dimensionality introduces significant challenges. To address this, we developed heuristics aimed at exploring and exploiting these spaces more effectively. These methods operate at a higher level of abstraction, focusing on identifying high-skew directions. Our findings demonstrate that these approaches, even though just heuristic, outperform traditional clustering methods in finding problematic regions within datasets.
@@ 10. Conclusion
In this paper, we introduced, formalized, and studied the problem of mining the unknown, underrepresented, and underperforming minority groups (Minoria). Making the connection to computational geometry concepts, we proposed efficient algorithms and practical solutions for finding high-skew projections as the combination of data features that reflect high-performance disparities at their tail. The scope of our proposed techniques is limited to the cases where the Minoria are linearly separable from the rest of the data. Extending our work to the more general class of non-linear projections, as well as developing alternative Minoria mining approaches, is an interesting direction for future work.